OpenStudy (anonymous):

help me finding L{tsin3t}

OpenStudy (anonymous):

may i ask what is L? is it just a constant? hope u could rewrite the question properly

hartnn (hartnn):

can u first find L{sin 3t} ??

OpenStudy (anonymous):

Laplace = L

hartnn (hartnn):

do u know L {sin 3t} ? @MadMike

OpenStudy (anonymous):

yea.!! \[a \div s ^{2} \times a^2\]

OpenStudy (anonymous):

srry not x its +

hartnn (hartnn):

\[\frac{a}{s^2+a^2}\]

OpenStudy (anonymous):

now tell me how to fine {t sin2t}

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

right ,so now there is a property,let,f(t)=sin3t,then F(s)=a/(s^2+a^2) L{t.f(t)}=-(derivative w.r.t s of F(s))

OpenStudy (anonymous):

so whts the derivative of 3/s^2 + 3^2

hartnn (hartnn):

u also need help in taking that derivative?

OpenStudy (anonymous):

yea i m noob

OpenStudy (anonymous):

please b quick

hartnn (hartnn):